OpenStudy (anonymous):

solve and check your answer. 17 over x+1 plus 3 over x+1=-4

OpenStudy (anonymous):

\[\frac{17}{x+1}+\frac{3}{x+1}=\frac{20}{x+1}\] is a start

OpenStudy (anonymous):

then \[\frac{20}{x+1}=-4\] \[20=-4(x+1)\] \[-5=x+1\] \[-6=x\]

OpenStudy (anonymous):

so x=-6?

OpenStudy (saifoo.khan):

Yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw